OpenStudy (anonymous):

a population of insects doubles every month. this particular population started out with 20 insects

OpenStudy (anonymous):

Umm youre missing some of his paragraph... How many months?

OpenStudy (mathstudent55):

Start with 20 After 1 month, 20 * 2 = 20 * 2^1 After 2 months, (20 * 2) * 2 = 20 * 2^2 After 3 months, (20 * 2^2) * 2 = 20 * 2^3 After x months: 20 * 2^x